OpenStudy (anonymous):

Medal/Fan <3 Which conjunction or disjunction equivalent to the open sentence 4 –3|n + 2| ≥ 1? A. n + 2 ≥ 1 and n + 2 ≤ –1 B. n + 2 ≤ 1 and n + 2 ≥ –1 C. n + 2 ≤ 1 and n − 2 ≥ −1 D. n + 2 ≥ 1 or n – 2 ≤ 1

OpenStudy (agent0smith):

First get the abs value alone, by subtracting 4 on both sides

OpenStudy (anonymous):

@amoodarya @jdoe0001 @pgpilot326 @ehuman

OpenStudy (amoodarya):

4 –3|n + 2| ≥ 1 –3|n + 2| ≥ 1-4 –3|n + 2| ≥ -3 divide by -3 |n + 2| <= 1

OpenStudy (amoodarya):

|a|<b mean -b<a<b can you go on ?

OpenStudy (anonymous):

okay @agent0smith ...

OpenStudy (anonymous):

Alright.

OpenStudy (agent0smith):

sorry, forgot a negative\[\large \left| ax + b \right| \le c \] \[\large -c \le ax + b \le c \]

OpenStudy (agent0smith):

You have \[\large \left| n+2 \right|\le 1\]so use the above

OpenStudy (anonymous):

Okay I get you. But I need an answer. Not just giving me the answer but I need to figure this out.

OpenStudy (agent0smith):

So try it using the last two posts. All you have to do is compare them.

OpenStudy (agent0smith):

\[\large \left| ax + b \right| \le c \]you have \[\large \left| n+2 \right|\le 1\] \[\large -c \le ax + b \le c\]so you'll have...

OpenStudy (anonymous):

do you understand why the sign switches directions when you multiply or divide both side of an inequality by a negative number?

OpenStudy (anonymous):

Nope... I'm finna get all F's that's why I'm studying.

OpenStudy (anonymous):

so let's make it easy and start with something we know is true, like \(-1 < 1\). this is true, yeah?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

now what happens when we multiply both sides by -1? what's on the left and what's on the right?

OpenStudy (anonymous):

umm this is where I get lost.

OpenStudy (anonymous):

what's -1 x -1?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

very good. and what's -1 x 1?

OpenStudy (anonymous):

-1 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now let's look at this all together... \[-1<1\]\[(-1)(-1)\overset{?}{<}(-1)(1)\]\[\Rightarrow 1\overset{?}{<}-1\]\[\Rightarrow 1>-1\] sign is switched whne we mulitply by a negative!

OpenStudy (anonymous):

Ahhhh okay.

OpenStudy (anonymous):

so from our original we have... \[4-3|n+2|\ge 1\] and subtracting 4 from both sides we get \[-3|n+2|\ge -3\] and now we divide both sides by -3 and switch the sign because we're dividing by a negative \[|n+2|\le 1\] and so we have isolated the absolute value! do you follow this?

OpenStudy (anonymous):

Okay I'm following but I'm still a few feet behind.

OpenStudy (anonymous):

take a minute and get there. i don't want to lose you. and if you have a question about a step or steps, please ask!

OpenStudy (anonymous):

Okay what exactly is "isolated" ???

OpenStudy (agent0smith):

Isolated means alone.

OpenStudy (anonymous):

isolated means by itself.

OpenStudy (anonymous):

Mkay cool so I got it

OpenStudy (anonymous):

do you see in the last step how |n+2| is isolated? nothing else on its side?

OpenStudy (anonymous):

Yes, sir.

OpenStudy (anonymous):

I'm assuming you're a sir.

OpenStudy (anonymous):

okay, so now remember the circle?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i'm a man, but not a sir. i havne't been knighted and didn't serve in the military.

OpenStudy (anonymous):

so draw the line, make the circle and remember what you're looking for.

OpenStudy (anonymous):

Oh okay... I live in Cali but my parents are from the south so I automatically assume all men are "sirs".

OpenStudy (anonymous):

 so to be inside the circle, n+2 must be less than 1 AND greater than -1. so to be on the circle, n+2 must be equal to 1 OR equal to -1. so to be outside the circle, n+2 must be greater than 1 OR less than -1.

OpenStudy (anonymous):

i live in "Cali" too and even though I'm in southern "Cali" it's not considered the south. silly civil war thing.

OpenStudy (anonymous):

haha I meant southeast. :)

OpenStudy (anonymous):

i know. is the problem making sense? the approach and solution?

OpenStudy (anonymous):

Yes it's making sense. :D And I'm in SoCal, too.

OpenStudy (anonymous):

lol, the south is being frozen right now and we're in the 70's... how lucky are we?

OpenStudy (anonymous):

VERY! My grandpa is sending me hate mail. He hopes everyone in SoCal gets worms or something.

OpenStudy (anonymous):

yuck! what a nice guy! sic

OpenStudy (anonymous):

Haha I still love him :) anyways, back on track.

OpenStudy (anonymous):

you good on this one?

OpenStudy (anonymous):

well I think so but is the answer D...? that runs in the family-understanding but getting the wrong answers. :/

OpenStudy (anonymous):

look back at what we did and then select the answer that matches what we came up with

OpenStudy (anonymous):

Oh duh.

OpenStudy (anonymous):

So I'm assuming (I assume too much) the answer is A?

OpenStudy (anonymous):